OpenStudy (anonymous):

Fool's problem of the day: There are two ants on opposite corners of a cube. On each move, they can travel along an edge to an adjacent vertex. If the probability that they both return to their starting position after \( 4 \) moves is \( \frac mn\) , where \( (m,n)=1 \), can you find \( m+n \)? (NOTE:They do not stop if they collide.) [The right answer was posted by @experimentX] My apologies I made the wrong assumption as @Arnab09 did. PS: If you Google it might spoil the problem for you.

OpenStudy (anonymous):

@nikvist, please check it out

OpenStudy (anonymous):

@callisto, what do u think?

OpenStudy (callisto):

Can they travel in the way like go forward by one side length, then go backward to the original position, then go forward by one side length, and then go backward to the original position?

OpenStudy (anonymous):

I guess so..

OpenStudy (callisto):

So, they have like 3 ways of travelling? 1. travel in a square form 2. forward -> forwarrd -> backward ->backward 3. forward -> backward -> forward -> backward

OpenStudy (anonymous):

square form.. there are 2 types

OpenStudy (callisto):

I meant like that..

OpenStudy (anonymous):

i got it..

OpenStudy (anonymous):

that square path can be in anti clockwise direction too..

OpenStudy (callisto):

but i think if i doesn't specify the first step, it doesn't matter?

OpenStudy (callisto):

Sight..* I don't, i need to do more grammar exercises :S

OpenStudy (callisto):

@experimentX I think so

OpenStudy (callisto):

*sigh

OpenStudy (anonymous):

well, i solved it this way: we assume the sides of the cube along the co-ordinate axes, one ant is at (0,0,0) and the other is at (1,1,1) from the beginning, each ant has 3 options to choose a path. after one move, the next move can be chosen is 3 ways. so, an ant can follow 3^4=81 different paths. to come back to home, within 4 moves, 2 are along chosen from +x, +y, +z other 2 are automatically selected.. they are among -x,-y,-z. we have to choose 2 moves from x, y, z.. this can be done in 3C2=6ways other moves are automatically selected.. now, among these 4 moves, we have to divide them in 2 groups.. for instance, from the point (0,0,0), positive moves have to come first, and from (1,1,1), negative moves have to come first.. so, there will be total 3C2*4!/(2!*2!)=36 moves possible for each ant to come back home.. so, probability to come back home for each ant= 36/81=4/9 as the moves of the ants are independent to each other, net probability to come back home for each ant= (4/9)^2=16/81 so, m+n= 16+81= 97 is it correct??

OpenStudy (anonymous):

sorry, i made a mistake earlier..

OpenStudy (experimentx):

ans is either 10 or 778

OpenStudy (experimentx):

it's 778

OpenStudy (anonymous):

Nobody has posted the right answer. Albeit @Arnab09 is very close! :)

OpenStudy (anonymous):

I wish I could do your problems, It would be fun being among you,trying to solve them , but I don't know , I'm dumb at these ! :( @FoolForMath

OpenStudy (anonymous):

i cant find whats wrong in my method :| if the ants collide, there path change?

OpenStudy (anonymous):

Is it safe to assume ants won't travel back to where they came from unless they collide with eachother?

OpenStudy (anonymous):

what a silly mistake by me!!! 3 C 2= 3

OpenStudy (anonymous):

i am making it correct again..

OpenStudy (anonymous):

well, i solved it this way: we assume the sides of the cube along the co-ordinate axes, one ant is at (0,0,0) and the other is at (1,1,1) from the beginning, each ant has 3 options to choose a path. after one move, the next move can be chosen is 3 ways. so, an ant can follow 3^4=81 different paths. to come back to home, within 4 moves, 2 are along chosen from +x, +y, +z other 2 are automatically selected.. they are among -x,-y,-z. we have to choose 2 moves from x, y, z.. this can be done in 3C2=3ways other moves are automatically selected.. now, among these 4 moves, we have to divide them in 2 groups.. for instance, from the point (0,0,0), positive moves have to come first, and from (1,1,1), negative moves have to come first.. so, there will be total 3C2*4!/(2!*2!)=18 moves possible for each ant to come back home.. so, probability to come back home for each ant= 18/81=2/9 as the moves of the ants are independent to each other, net probability to come back home for each ant= (2/9)^2=4/81 so, m+n= 4+81= 85 is it correct??

OpenStudy (anonymous):

@Foolformath, check this out..

OpenStudy (anonymous):

I think Arnab's right.

OpenStudy (anonymous):

thanx :D, @Foolformath

OpenStudy (anonymous):

Did you Googled? ;)

OpenStudy (anonymous):

nope.. honestly why? is it solved somewhere in same process..??

OpenStudy (anonymous):

Anyways good for you! :)

OpenStudy (anonymous):

where is the solution? its only a question sheet :O

OpenStudy (anonymous):

and if i would google it.. i wouldnt do mistake in 1st chance :)

OpenStudy (anonymous):

solution is not here

OpenStudy (anonymous):

variant=??

OpenStudy (anonymous):

There is some problem about my understanding with about problem I will address this issue sometime later.

OpenStudy (anonymous):

It's fixed now. Congrats to @experimentX!!

OpenStudy (anonymous):

post the solution link please?

OpenStudy (anonymous):

oh.. ok.. i got it now.. congrats to @experimentx

OpenStudy (experimentx):

sorry people, i couldn't find a better solution than to determine by counting the possible paths. it's was 21